OpenStudy (anonymous):

Describe the graph of a system of linear equations that has infinite solutions.

OpenStudy (anonymous):

Two lines right on top of each other.

OpenStudy (anonymous):

Same equations have infinite sollutions

OpenStudy (anonymous):

Like:- \[2x+4x = 8\] \[4x+8x=16\]

OpenStudy (sidsiddhartha):

Consider the two lines below y = 2x+1 and 2y = 4x +2. These two equations are really the same line . then they will have infinitr solutions and there graph will stay right on top of each other. like